OpenStudy (anonymous):

HELP THIS WORD PROBLEM SEEMS EASIER THAN YOU THINK. A dish of nuts contains cashews and peanuts. 2 more ounces of peanuts are added to the dish, making the new mixture 20% cashews. 2 more ounces of cashews are added, making the dish 33 1/3 % cashews. What percent of the original mixture of nuts was cashews?

OpenStudy (anonymous):

you have to know how many ounces of nuts are in the dish to begin with to solve the problem

OpenStudy (amistre64):

you have added the same ratio of pnuts and cashews so their ratio remains the same ...

OpenStudy (anonymous):

meaning...?

OpenStudy (amistre64):

if things remain the same ,,, then what does that tell us about the start and ending values?

OpenStudy (anonymous):

Then the start and end values would be of the same ratio to the other percents

OpenStudy (anonymous):

Amistre64 is incorrect, the ratio doesn't necessarily remain the same

OpenStudy (anonymous):

it would only remain the same if the starting ratio is 1:1

OpenStudy (amistre64):

are we assuming count or weight?

OpenStudy (anonymous):

do you know how to setup a system of equations?

OpenStudy (amistre64):

the only thing we have to go on is weight

OpenStudy (amistre64):

im open to being mistaken :)

OpenStudy (anonymous):

so am I :) so excuse me if I'm incorrect but I believe to solve this you need to setup a basic system of equations by modeling the two parts of the problem. Btw since this is my first time doing this, how much help should we give?

OpenStudy (amistre64):

hmm, guide, but let the asker participate. and correct as needed.

OpenStudy (anonymous):

Ahhh thank you! So Vincent, do you have an idea of how to do this problem?

OpenStudy (mathstudent55):

The dish started with p ounces of peanuts and c ounces of cashews. The original ratio is p/c and is unknown. Then 2 ounces of peanuts are added. Now the ratio is: (p + 2)/c = 4/1 Then, 2 ounces of cashews are added. The new ratio is (p + 2)/(c + 2) = 2/1

OpenStudy (anonymous):

Well... Not really. Ratios with different units confuse me a lot

OpenStudy (mathstudent55):

Solve the system of equations: \(\dfrac{p+2}{c} = 4\) \(\dfrac{p+2}{c+2} = 2\)

OpenStudy (anonymous):

so solve for c? Since we're looking for the cashew percent?

OpenStudy (anonymous):

mathstudent your equations are incorrect and not that helpful if he doesn't understand what he is doing.

OpenStudy (mathstudent55):

Find p and c, then find what percent of the total c is.

OpenStudy (anonymous):

1DEA, how's mathstudent55 incorrect?

OpenStudy (anonymous):

Oh wait actually they are right but do you understand how he is modeling those equations?

OpenStudy (anonymous):

Yea. Ratios of the cashews and ounces. I just solved both... C = 2

OpenStudy (mathstudent55):

If you don't understand how I arrived at the equations, read my explanation above.

OpenStudy (anonymous):

ok cool that's good if you understand how to do it, as it is more important to be able to arrive at those equations than just simply being able to solve the equations

OpenStudy (amistre64):

i was considering the rations as c --- = original percentage of c to (c+p) c+p c ------ = .2 c+p+2 c+2 ----- = .3333 c+p+4

OpenStudy (anonymous):

@mathstudent55, I briefly glanced at it initially and was confused for a sec.

OpenStudy (anonymous):

mathstudent, so if c= 2 , there was originally 2 ounces of cashews??

OpenStudy (mathstudent55):

@1DEA No problem.

OpenStudy (anonymous):

@amistre64 that's completely correct

OpenStudy (mathstudent55):

Yes, there were originally 2 ounces of cashews. Now you need the original number of ounces of peanuts.

OpenStudy (anonymous):

there were 6 ounces of peanuts

OpenStudy (anonymous):

so if there were 8 ounces in total... I just need to find 2 percent out of 8??

OpenStudy (mathstudent55):

yes

OpenStudy (mathstudent55):

2 is what percent of 8?

OpenStudy (anonymous):

25%

OpenStudy (mathstudent55):

correct

OpenStudy (anonymous):

Awesome. Thanks all of you!

OpenStudy (mathstudent55):

yw

OpenStudy (anonymous):

np